A.) First sample has n=3 and M=8. Second sample has n=5 and M=6. What is the weighted mean?

B.) First sample has n=10 and M=12. Second sample has n=4 and M=7. What is the weighted mean?


Also do you mind showing me the formula how to find the weighted mean?


1
Expert's answer
2022-02-01T13:49:54-0500

"M=\frac{M_1n_1+M_2n_2}{n_1+n_2}."


A) "M=\frac{8*3+6*5}{3+5}=\frac{54}{8}=6.75."


B) "M=\frac{12*10+7*4}{10+4}=\frac{148}{14}\approx 10.57."